### v1

# CoT-Space: A Theoretical Framework for Internal Slow-Thinking via Reinforcement Learning

Abstract
Reinforcement Learning (RL) has become a pivotal approach for enhancing the reasoning capabilities of Large Language Models (LLMs). However, a significant theoretical gap persists, as traditional token-level RL frameworks fail to align with the reasoning-level nature of complex, multi-step thought processes like Chain-of-Thought (CoT). To address this challenge, we introduce CoT-Space, a novel theoretical framework that recasts LLM reasoning from a discrete token-prediction task to an optimization process within a continuous, reasoning-level semantic space. By analyzing this process from both a noise perspective and a risk perspective, we demonstrate that the convergence to an optimal CoT length is a natural consequence of the fundamental trade-off between underfitting and overfitting. Furthermore, extensive experiments provide strong empirical validation for our theoretical findings. Our framework not only provides a coherent explanation for empirical phenomena such as overthinking but also offers a solid theoretical foundation to guide the future development of more effective and generalizable reasoning agents.
1 Introduction
Reasoning is a critical capability for Large Language Models (LLMs), and there has been a significant focus on developing methods to improve it. One of the most well-known frameworks is Chain-of-Thought (CoT) [1]. Under the CoT framework, a promising approach to enhance LLM reasoning is “test-time scaling”, also known as “slow thinking” [2, 3]. This branch of methods generally falls into two categories: external and internal. External methods extend the LLM’s output by introducing additional steps, often through a tree-structured search [4]. These techniques, which do not require extra training [5], include popular frameworks such as Monte-Carlo Tree Search (MCTS)[6, 7] and self-consistency (also known as Best-of-N) [8]. In contrast, internal methods integrate the slow-thinking capability directly into the model itself through post-training. This is achieved either by supervised fine-tuning (SFT), a process often referred to as distillation, or by reinforcement learning (RL).
Recent advancements in reasoning models such as DeepSeek’s R1 [9], OpenAI’s o1 and o3 [10, 11], and Qwen’s QwQ [12] have intensified the focus on the internal slow thinking methods, and especially using RL to enhance LLM reasoning. This training paradigm is often termed zero, indicating that the model is post-trained directly from its pre-training checkpoint using RL, bypassing the SFT stage. The results of zero-like approach are encouraging, as they demonstrate the potential for LLMs to develop reasoning capabilities through self-exploration with minimal supervision. This observation aligns with the core philosophy of RL: by setting a high-level, global objective, the policy can be guided to discover an optimal solution independently. As shown in Figure 1, the application of RL to LLM reasoning shares similarities with classical RL scenarios. For instance, as shown in subgraph (a), in a classic RL task like Bricks Breaker [^1] Bricks Breaker is a video game, the goal is to destroy the bricks by shooting a ball at them., the agent’s goal is to break more bricks. Without explicit instructions on specific strategies, the agent may discover a highly effective method, such as hitting the ball behind the top row of bricks to achieve a higher score. This kind of emergent behavior is a common outcome in classical RL. Similarly, in the context of LLM reasoning, as shown in subgraph (b), the policy model is given the broad goal of providing a correct answer. As RL training progresses, the model converges on an effective CoT pattern, demonstrating its ability to learn complex reasoning pathways from a simple, high-level reward signal.
Building on the discovery that models can automatically converge on a suitable CoT, various RL-based methods have been proposed for more effective zero-like post-training, including GRPO [13], VC-PPO [14], DAPO [15], and VAPO [16]. However, despite these practical advances, a significant gap remains in the theoretical understanding of LLM internal “slow thinking” via RL. Firstly, the surprising “overthinking” phenomenon [17] suggests there is an optimal CoT length for the reasoning process. Although empirical observations show different convergence patterns with various RL recipes, the precise mechanism driving this behavior remains unknown. Secondly, the generalization ability of LLM reasoning tasks is not yet well understood; while we observe this phenomenon, the underlying principles are not fully analyzed.
A central challenge to the theoretical analysis of LLM reasoning via RL is the misalignment between the traditional RL theoretical framework and the CoT paradigm. The conventional framework for RL with LLMs operates at the token level, where each action is the selection and concatenation of a single token to the current output. In contrast, the CoT approach functions as a higher-level reasoning framework, where the output is generated step-by-step, with each thought containing multiple tokens to solve a specific reasoning step. To address this limitation, we propose CoT-Space, a novel reasoning-level theoretical framework for LLM reasoning via RL. This framework is designed to enable a more precise theoretical analysis of this domain. We further validate its effectiveness by performing two corresponding analyses aimed at understanding both the convergence of CoT length and the generalization of reasoning tasks, thereby addressing the key theoretical gaps previously identified.
The remainder of this paper is structured as follows. In Section 2, we provide a systematic analysis of the misalignment between the current token-level RL formulation and the reasoning-level CoT framework. Subsequently, we introduce our reasoning-level theoretical framework, CoT-Space. We then leverage this framework in Section 3 to analyze the convergence of CoT length in the reasoning process from the perspective of noise and risk, demonstrating its potential theoretical value. We perform empirical validation of our theoretical insights in Section 4. Following this, we briefly review related works in Section 5 and finally conclude the paper in Section 6.
Figure 1: An illustration of the analogy between strategy discovery in classical RL and RL for LLM reasoning. (a) In a classical RL scenario such as a Brick Breaker game, the agent is given a high-level goal to break more bricks. Through exploration, it discovers a non-trivial, effective strategy to hit the ball behind the bricks to maximize its reward. (b) In LLM reasoning, the policy model is given a high-level goal to provide the correct answers. As RL training progresses, the model autonomously discovers that generating a suitable CoT is an effective strategy to achieve the goal.
2 CoT-Space: A Reasoning-Level Theoretical Framework
This section presents a systematic analysis of the limitations inherent in the current token-level RL formulation when applied to the higher-order cognitive processes of CoT reasoning. We first discuss the fundamental misalignment of this conventional approach in Subsection 2.1 and then discuss the continuum nature of the reasoning-level state space in Subsection 2.2. Finally, we introduce our novel, reasoning-level theoretical framework CoT-Space in Subsection 2.3.
2.1 From Token-Level to Reasoning-Level Perspective
The conventional theoretical framework for applying RL to LLMs is fundamentally misaligned with the reasoning-level nature of CoT scenarios. This token-level assumption is inherited from traditional RL frameworks, where the agent receives a reliable and meaningful reward at each state transition. For instance, classic RL tasks like chess, Go, or video games allow for a meaningful measurement of a state’s value, which guides the agent’s policy.
However, in the context of natural language, the auto-regressive nature of modern language models has led us to formulate each token prediction as a state transition, resulting in a popular token-level theoretical framework. We briefly review this formulation here. Prior works usually model natural language generation as a token-level Markov Decision Process (MDP): $\mathcal{M}(\mathcal{S},\mathcal{A},r,H)$, where $S$ is the state space of existing token sequences, $A$ is the action space representing the token dictionary, $r$ is the reward function guiding generation, and $H$ is the token budget. A policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ produces the next token, and the state transitions by concatenating the predicted token $a=\pi(s)$ to the existing sequence $s$. This formulation presents a significant challenge regarding the state space $\mathcal{S}$ as the following assertion.
Assertion 2.1.
The token-level state space $\mathcal{S}$ of natural language exhibits both exponential growth and discrete topology.
On the one hand, the state space $\mathcal{S}$ can be partitioned into a union of infinite subspaces based on sequence length $t$: $\mathcal{S}=\mathcal{S}_{0}\cup\mathcal{S}_{1}\cup\cdots\cup\mathcal{S}_{t}\cup\cdots$, where $\mathcal{S}_{t}=\mathcal{A}^{t}$. As the combinations of dictionary tokens grow exponentially, the subspace for a given prompt $q$, $\mathcal{S}_{|q|}$, and all subsequent subspaces $\mathcal{S}_{>|q|}$ are astronomically large. Consequently, the data observed during RL training constitutes only a minute fraction of the entire huge possible token-level state space. On the other hand, the transition between states is unidirectional and discrete due to the auto-regressive nature of the process: $\forall\bm{s}_{t}\in\mathcal{S}_{t},\mathcal{T}(\bm{s}_{t},\pi)\in\mathcal{S}_{t+1}$. This indicates the discrete nature of the token-level state space.
Turning to CoT reasoning processes, we observe that a complete reasoning trajectory of length $L$ comprises multiple conceptual steps $(\xi_{1},\xi_{2},\dots,\xi_{L})$ with each step $\xi_{l}$ realized through multiple token generations. Consider solving algebraic equations: each $\xi_{l}$ corresponds to a specific subgoal in the solution process, achieved through coordinated token generation. This perspective reveals two fundamentally distinct procedural layers:
  1. 1.
    Reasoning-Level: The strategic planning of correct reasoning steps to solve problems
  2. 2.
    Token-Level: The tactical execution of individual reasoning steps through token generation
Traditional MDP formulations reduce LLM reasoning to token-level exploration, while the essence of reasoning involves higher-order cognitive processes that orchestrate sequences of conceptual steps. The token-level perspective fails to capture the strategic planning and structural coherence inherent in human-like reasoning, creating substantial theoretical barriers to understanding modern reasoning paradigms such as CoT.
Thus, to establish theoretically sound foundations, we must transition from token-level formulations to reasoning-level perspectives that properly reflect the problem’s intrinsic nature.
2.2 On the Continuum Nature of the Reasoning-Level State Space
Building upon Assertion 2.1, which characterizes the token-level state space as both high-dimensional and discrete, we now provide a theoretical basis for approximating this space as a continuum at the reasoning level. The core insight is that while the token-level space consists of discrete points, the semantic space it represents becomes increasingly dense as the reasoning process unfolds.
The essence of reasoning lies in the semantic meaning of conceptual steps, not in the specific tokens used to express them. A single reasoning step, such as “calculating the total weekly pages” of the instance in Fig. 1, can be realized by a vast number of distinct token sequences:
  • “Step 2: He writes 6*2=12 pages every week.”
  • “Then, I’ll calculate the total weekly pages: 6 pages/friend * 2 friends = 12 pages.”
  • “So, for both friends, the page count per week is 6 $\times$ 2, which equals 12.”
These sequences correspond to distant, discrete states in the token-level space. However, in a higher-dimensional semantic space, they are nearly identical, occupying a very small neighborhood. The CoT-Space framework posits that as the CoT length $L$ increases, the number of semantically meaningful states grows so rapidly that the space they inhabit can be treated as a continuous manifold. This density arises from the expressive redundancy of language: the number of valid token-level paths to achieve a reasoning goal grows exponentially with the available token budget.
To formalize this, we introduce the following definitions and a key assumption.
Definition 2.2.
(Semantic equivalence set.) For an abstract reasoning step $\xi_{l}$ (e.g., a calculation, a logical deduction), its semantic equivalence set $\mathcal{T}(\xi_{l})$ is the collection of all token sequences $\tau\in\mathcal{A}^{*}$ that are semantically equivalent to $\xi_{l}$.
$\mathcal{T}(\xi_{l})=\{\tau\in\mathcal{A}^{*}\ |\ \text{Decode}(\tau)\equiv\xi_{l}\},$
where $\mathcal{A}^{*}$ is the space of all possible token sequences and $\text{Decode}(\cdot)$ is an abstract function mapping a token sequence to its semantic meaning.
Definition 2.3.
(Number of realizations.) For a complete reasoning trajectory of $L$ steps, $(\xi_{1},\xi_{2},...,\xi_{L})$, its total number of token-level realizations, $N_{\text{realize}}$, is the size of the Cartesian product of the semantic equivalence sets for each step.
$N_{\text{realize}}(\xi_{1},...,\xi_{L})=\prod_{l=1}^{L}|\mathcal{T}(\xi_{l})|.$
Definition 2.4.
(Reasoning state density.) Within a fixed semantic volume $\mathcal{V}_{\text{semantic}}$, the set of all valid, reachable reasoning states of reasoning token amount $K$ is denoted by $\mathcal{S}_{\text{reasoning}}^{(K)}$. The state density $\rho(L)$ is defined as the number of such states per unit of semantic volume.
$\rho(K)=\frac{|\mathcal{S}_{\text{reasoning}}^{(K)}|}{|\mathcal{V}_{\text{semantic}}|}.$
To substantiate our claim that this density grows exponentially, we propose the following intuitive assumption.
Assumption 2.5.
(Exponential expressive redundancy of language.) For any non-trivial reasoning step $\xi_{l}$, the size of its semantic equivalence set $|\mathcal{T}(\xi_{l},k)|$, when realized using an average of $k$ tokens, grows exponentially with $k$. That is, there exists a constant $c>1$ such that:
$|\mathcal{T}(\xi_{l},k)|\geq c^{k}.$
This assumption is grounded in the combinatorial nature of language; allowing even a few extra tokens for modifiers, rephrasing, or synonyms creates an exponentially larger space of semantically equivalent expressions.
Based on these foundations, we can now formally state the theorem that justifies the continuum approximation of reasoning-level state space.
Theorem 2.6.
(Continuum convergence of reasoning-level state space.) Under Assumption 2.5, the reasoning-level state density, denoted as $\rho(K)$, increases exponentially with the total reasoning token amount $K$. This exponential growth implies that the expected semantic distance between any two adjacent valid reasoning states decreases dramatically as $K$ increases, and approaches zero as $K\to\infty$. Therefore, in the macroscopic limit, the discrete reasoning-level state space converges to a continuous semantic manifold.
The proof is provided in Appendix A.1. Theorem 2.6 provides the formal justification for the central premise of our work. This result licenses our departure from the discrete, token-level MDP formulation and enables the adoption of more powerful, analytically tractable tools from continuous optimization theory. With the continuum nature of the reasoning space established, we are now positioned to formally introduce our framework. In the subsequent section, we will leverage this continuum perspective to define the CoT-Space and reframe the LLM reasoning process as an optimization problem within this newly characterized landscape.
2.3 Formulation of CoT-Space
Building upon the discussion above, which characterizes the token-level state space and illustrates the continuum nature of reasoning-level state space, we now develop a reasoning-level theoretical framework designed to bridge the analytical gap. This alternative perspective draws an analogy to interstellar travel in space, as shown in Figure 2. From a token-level perspective, each individual token and its corresponding state are like discrete points, or stars, in a vast semantic galaxy. The auto-regressive generation process then resembles a journey across these stars. Since transitions only occur between these discrete states, the generation process is fundamentally discrete at the token-level scale, as illustrated in panel (a).
However, from the reasoning-level perspective, a continuum approximation emerges when we “zoom out” to consider the limiting behavior of this discrete system. Due to the exponential density of semantically valid states, the global properties of the space converge toward continuum-like characteristics, as visualized in panel (b). This macroscopic viewpoint allows us to move beyond individual tokens and analyze reasoning from a more abstract, strategic level.
Figure 2: An illustration for the shift from a token-level to a reasoning-level perspective. (a) The traditional token-level scale views the generation process as a sequence of discrete transitions in a vast state space. (b) The proposed reasoning-level scale zooms out to view the state space from a macroscopic perspective, where the incredibly dense and numerous token-level states can be approximated as a continuous semantic reasoning space.
To enable this higher-level analysis, we provide the following definitions for a new formalization of LLM reasoning via RL. We begin by defining the reasoning-level state. Based on the CoT setting, where the output is generated step-by-step and consists of multiple intermediate thoughts, a reasoning state is defined as follows.
Definition 2.7.
(CoT reasoning state.) Given a query $q$ and an incomplete set of intermediate reasoning steps $\xi_{o}$, the current reasoning state is defined as $s_{o}=(q,\xi_{o})$.
Definition 2.7 establishes a reasoning-level state that is a more meaningful unit of analysis than a single token. We then introduce a special class of states that represent successful problem solutions, which we call minimums.
Definition 2.8.
(Minimums in reasoning space.) Given a query $q$ with a definite golden answer $\phi$. Assume $\Xi_{q}$ is the set of all reasonable intermediate reasoning process for query $q$. Then the minimums of the reasoning space are defined as $\bm{M}_{q}=\left\{(q,\bm{\xi},\phi)~|~\bm{\xi}\in\Xi_{q}\right\}$.
According to Definition 2.8, minimums are special states where the query has been successfully solved. For any intermediate reasoning state, we can then define which minimums are reachable from it.
Definition 2.9.
(Reachable minimums.) Given a query $q$ with a definite golden answer $\phi$, and an incomplete set of intermediate reasoning steps $\xi_{o}=[\xi_{1},\xi_{2},\dots,\xi_{t}]$. The reachable minimums are defined as $\bm{R}^{o}_{q}=\left\{(q,\bm{\xi},\phi)~|~\xi_{o}\sqsubseteq\bm{\xi}~\text{and}~\bm{\xi}\in\Xi_{q}\right\}$, where $a\sqsubseteq b$ represents that $a$ is prefix of $b$.
Definition 2.9 established the concept of reachable minimums. This concept represents the set of successful reasoning paths that can be completed from the current state. Building on this, we quantify the distance between a state and its reachable minimums and define the nearest minimum.
Definition 2.10.
(Reachable distance and nearest reachable minimum.) Given a query $q$ with a definite golden answer $\phi$, and a reasoning state $s_{o}$ on this query, the reachable distance between the $s_{o}$ and a reachable minimum is defined as $\operatorname{dist}(s_{o},m)=|\bm{\xi}|-|\xi_{o}|$, where $m\in\bm{R}^{o}_{q}$, $\bm{\xi}$ and $\xi_{o}$ are intermediate steps of $m$ and $s_{o}$ respectively. The nearest minimum is defined as $m^{*}_{o}=\underset{m\in\bm{R}^{o}_{q}}{\arg\min}\,\operatorname{dist}(s_{o},m)$.
Definition 2.10 quantifies the number of reasoning steps required to reach the closest successful solution from a given state. Finally, we define a reasoning loss to characterize the gap between the current state and the optimal solution.
Definition 2.11.
(Reasoning loss.) Given a query $q$ with a definite golden answer $\phi$, and a reasoning state $s_{o}$ on this query, there exists a measure $C$ that quantifies the loss of $s_{o}$ towards the golden answer, which satisfies: for two different states $s_{i},s_{j}$, $C(s_{i})<C(s_{j})$ if and only if $\operatorname{dist}(s_{i},m^{*}_{i})<\operatorname{dist}(s_{j},m^{*}_{j})$. Specially, $\forall m\in\bm{M}_{q},C(m)=0$.
Definition 2.11 conceptualizes the reasoning process as a descent in a loss landscape. The function $C(\cdot)$ acts as a proxy for the distance to a correct solution, where a lower loss indicates a more advanced stage of reasoning. For instance, consider the problem in Figure 1. A state $s_{i}$ with the intermediate thought “He writes $12$ pages every week” is closer to the final answer than a state $s_{j}$ with only “He writes each friend $32=6$ pages a week.” According to this definition, because $s_{i}$ requires fewer steps to reach a minimum (i.e., $\operatorname{dist}(s_{i},m_{i}^{*})<\operatorname{dist}(s_{j},m_{j}^{*})$), its reasoning loss must be lower, satisfying $C(s_{i})<C(s_{j})$.
With this complete reasoning-level perspective, we can now model the LLM reasoning process as an optimization problem rather than a token-level MDP. This formulation simplifies the analytical foundation by reframing the process as an update to the reasoning state, guiding it toward one of the solution minima.
3 Analytical Applications for Overthinking within CoT-Space
This section demonstrates the analytical utility of the CoT-Space framework by applying it to a critical issue in LLM reasoning: the existence of an optimal thought process length and the related problem of “overthinking”. Moving beyond traditional analyses, we utilize the continuous perspective of CoT-Space to provide a theoretical foundation for this empirical phenomenon. Our analysis proceeds in two parts: we first examine the reasoning process through the lens of optimization and noise (Subsection 3.1), and second, we assess its generalization capabilities from the perspective of risk (Subsection 3.2). Finally, we briefly conclude our theoretical insights in Subsection 3.3. This application of CoT-Space showcases its potential to ground empirical findings in a robust theoretical model.
3.1 Reasoning as Optimization: A Perspective of Noise
Figure 3: An illustration of LLM reasoning modeled as an optimization process. (a) In traditional ML, the goal is parameter optimization, where an algorithm iteratively calculates a gradient from data to update model parameters and converge to a minimum in the loss landscape. (b) In the CoT-Space framework, LLM reasoning is framed as state optimization, where the policy takes the current state and instruction to generate a new reasoning step, thereby updating the state to move closer to a solution minimum. (c) This panel visualizes the optimization process on a reasoning loss landscape, analogizing the CoT length to the learning rate in ML. A suitable CoT length, much like an optimal learning rate, navigates the space effectively to converge to a high-quality, generalizable solution.
By treating reasoning as an optimization process, we can draw a direct comparison to traditional machine learning (ML), as illustrated in Figure 3. As shown in subgraph (a), traditional ML optimization algorithms, such as SGD or Adam, formulate a cycle where the algorithm calculates a gradient based on the current parameters and training data to iteratively update the model. Similarly, in subgraph (b), when an LLM performs reasoning, its internal policy takes the current reasoning state and query instruction as input. It then generates a new step, which updates the state by concatenating the new output. This process establishes a parallel optimization cycle for LLM reasoning.
Such a comparison inspires an application of the sharpness-aware minimization theorem from classical ML. This theorem posits that an optimal amount of noise in the optimization process leads to a converged minimum with better generalization ability. As shown in subgraph (c) with the blue legends, the learning rate can introduce noise into the gradient descent process. A learning rate that is too large will fail to converge, while one that is too small may converge to a sharp, low-quality minimum. Only a suitable learning rate introduces the right amount of noise to converge to a flat, high-quality minimum. Given the similarities in the nature of optimization, this theoretical analysis can be applied to the reasoning process of LLMs, as shown with the orange legends in subgraph (c). In this analogy, the CoT length is analogous to the learning rate. A CoT that is too short may fail to produce a correct answer, while a CoT that is too long may lead to “overthinking”, resulting in a lower-quality response. Therefore, a suitable CoT length is essential for generating a high-quality response.
Building on this, we can model the optimization of LLM reasoning as a form of reasoning loss descent and perform a theoretical analysis. As previously established, when viewed from a reasoning-level perspective, the reasoning space can be considered approximately continuous, i.e., $s_{o}\in\mathcal{S}$ where $\mathcal{S}$ is a continuous space. The state is then optimized continuously based on the policy and reasoning loss. This perspective enables us to analyze the reasoning process through a loss-reduction lens.
For an $L$-step CoT reasoning process, this can be viewed as an $L$-step optimization based on the gradient of the reasoning loss. A single step of this optimization can be written as:
$\Delta s=-\frac{1}{L}\frac{\mathrm{d}}{\mathrm{d}s}\hat{C}$, (1)
where $\hat{C}$ is a noisy estimation of the reasoning loss, influenced by the policy and random sampling. We assume that $\hat{C}$ is a random variable with expectation $C$. Here, $L$ represents the total reasoning steps, or the CoT length.
Despite the above gradient descent formulation, this optimization process can also be modeled as a stochastic differential equation (SDE) with $L$ steps:
$\frac{\mathrm{d}s}{\mathrm{d}t}=-\frac{\mathrm{d}C}{\mathrm{d}s}+\eta(t),$ (2)
where $\eta(t)$ is the noise term and $t={1,\dots,L}$. We assume the noise term $\eta(t)$ is white noise with the statistical properties: $\langle\eta(t)\rangle=0$ and autocorrelation $\langle\eta(t)\eta(t^{\prime})\rangle=gF(s)\delta(t-t^{\prime})$, where $g$ is the noise scale, $F(s)$ s a function characterizing the magnitude of the reasoning loss estimation error at state $s$.
Equations (1) and (2) formulate the optimization process of CoT reasoning from the gradient update and SDE perspective, respectively. By combining the two equations related to noise, we can thus obtain the relationship between noise scale and CoT length, as shown in the following theorem.
Theorem 3.1.
(Relationship between noise scale and CoT length.) In a given optimization space, the noise scale $g$ is inversely proportional to the reasoning length $L$, i.e. $g\propto\frac{1}{L}$.
The proof is provided in Appendix A.2. Theorem 3.1 clarifies the relationship between the noise $g$ and the CoT length $L$ in the same the reasoning process. According to the sharpness-aware minimization theorems, given a reasoning task, there exists an optimal noise scale $g_{\text{opt}}$ in the reasoning process, which corresponds to an optimal CoT length $L_{\text{opt}}\propto\frac{1}{g_{\text{opt}}}$. In essence, this theoretical analysis provides a fundamental basis for why an optimal CoT length is crucial for effective reasoning, highlighting its role in introducing the right amount of noise to achieve a flat, high-quality, and generalizable solution.
3.2 Reasoning and Generalization: A Perspective of Risk
The previous section considered reasoning from the perspective of noise and analyzed it through an optimization lens. Here, we analyze the generalization risk of a policy to provide further theoretical insights, focusing on the upper bounds of reasoning risks.
In the context of CoT-Space, risks can be defined similarly to classical ML. The total expected error of a policy $\pi$ on the true data distribution $\mathcal{D}$, denoted by $R(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]$, can be decomposed into two distinct components:
$R(\pi)=\underbrace{\hat{R}_{n}(\pi)}_{\text{Empirical Loss (Bias)}}+\underbrace{\left(R(\pi)-\hat{R}_{n}(\pi)\right)}_{\text{Generalization Error (Variance)}}.$ (3)
Here, $\hat{R}_{n}(\pi)=\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi})$ is the empirical loss on the training set. In the following parts, we will analyze how each component behaves as a function of the CoT complexity, which we proxy by the expected CoT length $L$.
3.2.1 An Upper Bound for Generalization Error: The Overfitting Risk
We define the reasoning generalization error as the difference between the true expected loss and the empirical loss. This error quantifies a policy’s ability to generalize its reasoning to unseen queries, a concept parallel to the generalization error in traditional machine learning theory.
Definition 3.2.
(Reasoning generalization error.) Given $n$ queries $\{q_{i}\}_{i=1}^{n}$ sampled i.i.d. from a distribution $\mathcal{D}$, and a policy $\pi$ learned from $\{q_{i}\}_{i=1}^{n}$, let’s denote $s_{q}^{\pi}=\pi(q)$ as the end reasoning state output by $\pi$ on query $q$. The reasoning generalization error of $\pi$ on $\mathcal{Q}$ is defined as:
$R(\pi)-\hat{R}_{n}(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]-\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi}).$
To derive an upper bound for this error in Definition 3.2, we can simply adapt the PAC-Bayes framework, which provides a powerful link between generalization, policy complexity, and data volume. For our analysis within CoT-Space, we establish the following components:
  • Hypothesis Space $\mathcal{H}$: The space of all possible reasoning policies, denoted by $\mathcal{H}$. Our learned policy $\pi\in\mathcal{H}$.
  • Prior Distribution $P$: A distribution over the policy space $\mathcal{H}$ reflecting our initial beliefs about which policies are "good," often favoring simpler ones.
  • Posterior Distribution $Q$: A distribution over $\mathcal{H}$ that is learned from the training data queries $\{q_{i}\}_{i=1}^{n}$, concentrating on policies that perform well on that data.
By applying the PAC-Bayes theorem [18], we obtain a bound on the expected true loss which, when rearranged, yields an upper bound on the expected generalization error.
Theorem 3.3.
(PAC-Bayesian generalization upper bound for LLM reasoning.) With CoT-Space defined above, the policy $\pi$ is trained on an i.i.d. drawn training set with size $n$,by assuming the reasoning loss is bounded in $[0,1]$, for any prior $P$, with probability at least $1-\delta$, the following holds for all posterior $Q$:
$\mathbb{E}_{\pi\sim Q}[R(\pi)-\hat{R}_{n}(\pi)]\leq\sqrt{\frac{\operatorname{KL}(Q||P)+\ln\frac{2n}{\delta}}{2n}}$
where $\operatorname{KL}(Q||P)=\mathbb{E}_{\pi\sim Q}[\ln\frac{Q(\pi)}{P(\pi)}]$ is the Kullback-Leibler (KL) divergence between the posterior and prior distributions.
A detailed analysis is provided in Appendix A.3. Theorem 3.3 provides a powerful, high-probability upper bound on the expected reasoning generalization error for policies sampled from the learned posterior $Q$. While this PAC-Bayesian bound is a valid, generic result, we can derive a more specific bound by using information theory, a perspective that connects the generalization error to the mutual information between the learned policy and the training data. This approach leads to the following information-theoretic upper bound.
Theorem 3.4.
(Information-theoretical generalization upper bound for LLM reasoning.) With CoT-Space defined above, if the policy $\pi$ is trained on an i.i.d. drawn training set $S$ with size $n$, the following upper bound for the reasoning generalization error holds:
$|\mathbb{E}[R(\pi)-\hat{R}_{n}(\pi)]|\leq\sqrt{\frac{C_{max}^{2}\cdot\left(\mathbb{E}_{\pi,S}[L]\cdot\mathbb{E}_{\pi,S}[|\xi|]\right)\cdot\log|\mathcal{A}|}{2n}},$
where $C_{max}$ is the maximum possible value of the reasoning loss $C(\cdot)$, $L$ is the number of reasoning steps in a CoT data, $|\xi_{i}|$ is the number of tokens in the $i$-th reasoning step in a CoT data.
A brief proof is provided in Appendix A.4. Theorem 3.4 explicitly connects the generalization error to the observable, structural properties of the reasoning process. It suggests that overfitting, or “overthinking”, can occur in two distinct ways: through excessive reasoning depth (high $\mathbb{E}[L]$) or through excessive step verbosity (high $\mathbb{E}[|\xi|]$), both resulting in overlong response length. This provides a direct theoretical basis for regularizing the CoT generation process in RL to improve reasoning generalization.
3.2.2 A Lower Bound for Empirical Loss: The Underfitting Risk
The empirical loss, $\hat{R}_{n}(\pi)$, measures a policy’s ability to solve the problems it has already encountered. For complex reasoning tasks, a certain minimum number of steps is required for a correct solution. Intuitively, if $L$ is too small, the policy lacks the necessary “thinking space” to solve the queries, even those in the training set. This leads to high empirical loss and underfitting. As $L$ increases, the policy has more capacity to find the correct reasoning path, so the empirical loss $\hat{R}_{n}(\pi)$ generally decreases.
To formalize this, we quantify the notion of problem complexity and establish a condition for policy failure.
Definition 3.5.
(Required reasoning depth.) For any given query $q$, we define its Required Reasoning Depth, denoted as $L^{*}(q)$, as the minimum number of CoT steps required to reach any success state $m\in M_{q}$ from the initial state $(q,\emptyset)$. This value is an intrinsic property of the query’s complexity.
Definition 3.5 can be likened to the minimum number of steps required to solve a reasoning problem. For instance, a simple two-digit addition problem might require only one or two mental steps, making its $L^{*}$ very low. In contrast, proving a complex theorem, such as the Pythagorean theorem, requires a sequence of many interconnected logical steps, resulting in a high $L^{*}$. A policy that attempts to solve the theorem with only a few steps, regardless of how smart it is, will inevitably fail.
Building on the concept of $L^{*}$, we now formalize the direct consequence of a policy’s generated reasoning being too shallow. We make a straightforward assumption about the relationship between insufficient reasoning depth and policy failure, which translates a structural property of the CoT into a quantifiable error.
Assumption 3.6.
(Policy failure on insufficient depth) When a policy $\pi$ generates a CoT of length $L_{i}$ for a query $q_{i}$, if the length is insufficient to solve the problem, i.e., $L_{i}<L^{*}(q_{i})$, the policy is guaranteed to fail. The resulting reasoning loss $C(s_{q_{i}}^{\pi})$ is therefore bounded below by a significant positive constant, which we denote as $C_{fail}>0$. In the worst case, this loss can be as high as $C_{max}$.
Assumption 3.6 claims that a policy’s failure due to insufficient depth can be observed in practice. Consider a complex physics problem requiring multiple steps: identifying the forces, setting up equations, and solving for variables. A policy that only completes the first step and then stops will not be able to reach the correct final answer. In this case, the loss function would assign a high penalty (our $C_{fail}$) because the generated CoT, while possibly correct for the initial step, is fundamentally incomplete for solving the problem. The policy, therefore, cannot even fit the training data for problems that require deeper reasoning.
With this framework established by Definition 3.5 and Assumption 3.6, we can derive a lower bound for the empirical reasoning loss $\hat{R}_{n}(\pi)$.
Theorem 3.7.
With CoT-Space defined above, if the policy $\pi$ is trained on an i.i.d. drawn training set, the following lower bound for the empirical reasoning risk holds:
$\hat{R}_{n}(\pi)\geq P_{\pi}(L<L^{*})\cdot C_{fail},$
where $P_{\pi}(L<L^{*})=n_{fail}/n$ is the empirical failure rate of policy $\pi$ on the training set.
A detailed proof is provided in Appendix A.5. Theorem 3.7 provides a formal basis for the “underfitting” portion of our trade-off curve. It demonstrates that the empirical loss is directly constrained by the probability that the policy generates a CoT that is too short for the problem’s intrinsic complexity. A policy that tends to produce short CoTs (i.e., has a small $\mathbb{E}[L]$) will have a high failure rate $P_{\pi}(L<L^{*})$ on any dataset containing non-trivial problems. Consequently, its empirical loss $\hat{R}_{n}(\pi)$ is guaranteed to be high. This formalizes why a policy must learn to increase its CoT length $L$ to achieve a low empirical loss, thus completing our explanation for the tradeoff in our total error analysis and the existence of an optimal $L_{\text{opt}}$.
3.2.3 Tradeoff between Overfitting and Underfitting Risk
In Theorem 3.4, we provide an upper bound for the generalization error in the reasoning process with relation to CoT length $L$. While in Theorem 3.7, we provide a lower bound for the empirical loss in the reasoning process with relation to CoT length $L$.
Since the total expected error $R(\pi)$ is the sum of a decreasing function (empirical loss) and an increasing function (generalization error), this classic Bias-Variance Trade-off inevitably leads to a U-shaped curve for the total error as a function of the CoT length $L$. For small $L$, the total error is high due to high bias (underfitting). The policy can’t solve even the training problems. For large $L$, the total error is high due to high variance (overfitting). The policy has too much capacity and learns non-generalizable details.
Therefore, an optimal length, $L_{\text{opt}}$, must exist that minimizes the total error. This $L_{\text{opt}}$ is long enough to avoid underfitting but short enough to prevent significant overfitting. Our analysis reveals that there is no universal best CoT length, instead, the existence of an optimal length is a natural consequence of the fundamental trade-off between fitting the training data and generalizing to new data. From this perspective, any RL algorithm that can explore effectively can eventually converge to this optimal strategy.
3.3 Takeaways
We have analyzed with the CoT-Space from two perspectives: that of noise (Subsection 3.1) and risk (Subsection 3.2). Based on the theoretical insights from Theorems 3.1, 3.4, and 3.7, we can draw the following key conclusions about the factors that govern the convergence of the optimal CoT length $L_{\text{opt}}$ in LLM reasoning via RL.
Remark 1. The intrinsic difficulty of a reasoning task is a primary determinant of the optimal CoT length. According to Theorem 3.7, a more challenging task necessitates a greater $L^{*}$ to achieve a low empirical loss and avoid underfitting. Consequently, the optimal CoT length, $L_{\text{opt}}$, must be sufficiently large to accommodate the task’s complexity, otherwise the policy’s failure rate will remain high. This insight formalizes the intuitive idea that complex problems require more thinking steps to be solved correctly, and any effective policy must learn to generate a CoT of a length commensurate with the problem’s difficulty.
Remark 2. The capacity of the policy model directly influences the optimal CoT length by affecting the risk of overfitting. A model with higher capacity (e.g., more parameters) can explore a more expansive reasoning state space. This expanded exploration enables it to access states with higher reasoning losses that are inaccessible to weaker models, leading to a larger effective $C_{max}$. The generalization bound in Theorem 3.4 is an increasing function of both $C_{max}$ and the CoT length $L$. Consequently, to mitigate the heightened overfitting risk posed by its own high capacity and the resulting larger effective $C_{max}$, a more powerful model must learn to regularize its output. It achieves this by converging to a smaller optimal CoT length ($L_{\text{opt}}$), thereby maintaining a low generalization error.
Remark 3. While the choice of RL algorithm is crucial for effective training, our analysis suggests it has a limited direct influence on the final, converged CoT length. The optimal CoT length ($L_{\text{opt}}$) is primarily a function of the policy’s capacity and the inherent complexity of the training data. The RL algorithm’s main role is to find a policy that converges to this optimal length. As long as the algorithm is effective in its optimization, the final $L_{\text{opt}}$ is a property of the learned policy and the data distribution, rather than a byproduct of the specific RL approach used.
Remark 4. In a given problem space, the scale of noise in the optimization process is inversely correlated with the converged CoT length. According to Theorem 3.1, when a policy is exploring the same task space, a higher level of noise will lead to a smaller converged optimal CoT length ($L_{\text{opt}}$). This implies that a noisier optimization environment tends to favor simpler, shorter reasoning paths, likely because longer CoTs are more susceptible to the disruptive effects of noise. This finding underscores the importance of a stable training environment for achieving complex and in-depth reasoning.
To substantiate these theoretical derivations and demonstrate their practical relevance, we will now proceed to a series of empirical investigations in the following section. These experiments are designed to systematically test the four central remarks, providing robust empirical evidence for our proposed framework.
4 Experiments
This section presents a series of experiments designed to empirically validate the theoretical insights established in the preceding sections. After detailing the experimental setup in Subsection 4.1, we systematically investigate the relationship between the optimal CoT length ($L_{\text{opt}}$) and key factors, including task difficulty, model capacity, the choice of RL algorithm, and optimization noise scale in Subsections 4.2 through 4.5, respectively.
4.1 Experimental Settings
Dataset The experiments are conducted on the MATH [19] and GSM8K [20] benchmarks. The MATH training set is manually subdivided into five difficulty levels (Level 1–5) according to the original “level” field, where Level 5 is substantially more challenging than Level 1.
Base Model A diverse collection of pre-trained LLMs with varying architectures, parameter counts, and versions serves as the reference model for the reinforcement-learning algorithms. Specifically, the models include
  • Qwen2.5 series [21]: Qwen2.5-7B, Qwen2.5-3B, Qwen2.5-1.5B, and Qwen2.5-0.6B.
  • Qwen3 series [22]: Qwen3-8B-Base, Qwen3-4B-Base, Qwen3-1.7B-Base, and Qwen3-0.6B-Base.
  • Llama3 series [23]: Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, and Llama-3.2-1B-Instruct.
The maximum response length is set to 2,048 tokens for the Qwen2.5 and Llama3 models and to 8,196 tokens for the reasoning-intensive Qwen3 models.
Algorithm and Hyperparameters The evaluation covers the current mainstream reinforcement-learning algorithms for mathematical reasoning: GRPO [13], PPO [24], DAPO [15], Reinforce++ [25], and RLOO [26]. The reward function is rule-based [9]. For PPO, the critic model is initialized from Qwen2.5-0.6B. The learning rate for critic training is $1\times 10^{-5}$, the k3 KL-regularization coefficient is 0.001, and the GAE parameters $\lambda$ and $\gamma$ are both set to $1$. For DAPO, no KL regularization is applied; the high and low clip ratios are $0.28$ and $0.20$, respectively, and $L_{cache}$ is $1,024$. Across all algorithms, each training sample is sampled eight times ($K=8$). The actor is optimized with AdamW [27] at a learning rate of 1e-6, the training batch size is $128$, and training proceeds in a purely on-policy manner for up to 50 steps. All experiments are run on 8 $\times$ 96 GB NVIDIA H20 GPUs.
Figure 4: Empirical Validation of the Factors Governing Optimal CoT Length ($L_{\text{opt}}$). (a) The relationship between task difficulty and $L_{\text{opt}}$, demonstrating that more complex tasks require longer reasoning process. (b) The inverse relationship between model capacity and $L_{\text{opt}}$, where more powerful models converge to shorter CoTs to mitigate overfitting. (c) The convergence of mean response length under different RL algorithms, showing that the final $L_{\text{opt}}$ is largely algorithm-agnostic. (d) The inverse correlation between optimization noise scale (proxied by KL-Loss) and $L_{\text{opt}}$, indicates that noisier training environments favor more concise reasoning.
4.2 Task Difficulty and Converged Length
To empirically validate Remark 1, we investigate the relationship between task difficulty and the converged optimal CoT length ($L_{\text{opt}}$). In this experiment, we utilize the converged mean response length of the policy (Qwen2.5-7B-Base) as a direct empirical measure for $L_{\text{opt}}$. The intrinsic difficulty of the task is stratified by using datasets of increasing complexity, namely GSM8K and the five manually curated levels of the MATH dataset (Level 1 to Level 5). As illustrated in Figure 4(a), there is a clear and monotonic trend: as the complexity of the task increases, the corresponding $L_{\text{opt}}$ also increases significantly. This result provides strong evidence for our theoretical claim. According to Theorem 3.7, more challenging tasks possess a greater required reasoning depth ($L^{*}$). Consequently, the policy must generate a longer CoT to avoid the high empirical loss associated with underfitting. The experimental outcome confirms that the RL process successfully identifies and converges to a longer $L_{\text{opt}}$ to accommodate the inherent complexity of more difficult problems. We also provide validations on more base models in Appendix B.
4.3 Model Capacity and Converged Length
This subsection examines the influence of model capacity on the optimal CoT length, thereby verifying Remark 2. We train the models by GRPO, then use the parameter count of different models as a proxy for their capacity and the final critic score as a measure of their reasoning performance. Figure 4(b) plots the converged CoT length ($L_{\text{opt}}$) against this critic score for various models trained on MATH. The results reveal a distinct negative correlation: within each model family (e.g., Llama 3, Qwen 2.5, and Qwen 3), larger and higher-performing models consistently converge to a shorter $L_{\text{opt}}$. This empirical finding aligns precisely with our theoretical analysis. As stated in Remark 2 and derived from the generalization bound in Theorem 3.4, models with higher capacity are inherently more susceptible to overfitting. The risk of overfitting is amplified by longer and more complex CoTs. Therefore, to minimize the total reasoning error, a more powerful model learns to regularize its reasoning process by adopting a more concise, shorter CoT, effectively balancing the trade-off between empirical performance and generalization.
4.4 RL Algorithm and Converged Length
In this experiment, we assess the impact of different RL algorithms on the final converged CoT length to test the validity of Remark 3. We track the mean response length over the course of training for five different RL algorithms: REINFORCE++, RLOO, PPO, DAPO, and GRPO. The resulting dynamics are displayed in Figure 4(c). While the convergence trajectories and learning speeds differ markedly among the algorithms—for example, GRPO gradually increases the response length, whereas RLOO maintains a high length from the outset—all tested algorithms ultimately steer the policy toward a similar range of converged mean response lengths for a given task and model. This outcome supports our insight in Remark 3 that the choice of RL algorithm is primarily a facilitator of the optimization process rather than a determinant of the optimal solution’s properties. The final converged length, $L_{\text{opt}}$, appears to be an intrinsic characteristic dictated by the interplay between model capacity and task difficulty, not an artifact of the specific optimization algorithm used to find it.
4.5 Noise Scale and Converged Length
Finally, we validate Remark 4 by analyzing the relationship between the converged CoT length and the noise scale of the optimization process. For this analysis, we employ the Kullback-Leibler (KL) divergence loss, or “KL-Loss”, as a quantifiable proxy for optimization noise. A higher KL-Loss indicates a greater deviation from the reference policy, representing a less stable and thus “noisier” update. Figure 4(d) illustrates the relationship between the final KL-Loss and the converged length ($L_{\text{opt}}$) across various models and task difficulties. The data shows a clear inverse correlation: for a given task, models that exhibit a higher KL-Loss consistently converge to a shorter $L_{\text{opt}}$. This empirical result substantiates the relationship derived in Theorem 3.1. A noisier optimization environment penalizes longer, more intricate reasoning paths, as they are more susceptible to disruption from unstable policy updates. Consequently, the RL process naturally favors shorter and more robust CoTs to ensure stable convergence, confirming that a higher noise scale leads to a smaller optimal CoT length.
5 Related Work
5.1 Reasoning with LLMs
Recent advancements in Large Language Models (LLMs) have significantly enhanced their capacity for complex reasoning. Building upon the foundational concept of LLMs as few-shot learners [28], a pivotal development has been the use of prompting techniques to elicit explicit reasoning processes. Chain-of-Thought (CoT) prompting [1], for instance, guides models to generate intermediate steps, a method whose robustness was later improved by Self-Consistency through the aggregation of multiple reasoning paths [8]. Theoretical understanding has advanced in parallel, with information-theoretic frameworks offering methods to quantify and mitigate reasoning errors, thus bridging empirical and theoretical work [29]. Concurrently, explorations into the representation of reasoning have yielded novel formats, from structured code in Program-of-Thought [30] to internal thought tokens generated within the model’s latent space [31, 32]. However, these explorations have been largely empirical, often lacking a strong theoretical foundation.
5.2 RL for LLM Reasoning
Test-time scaling has emerged as an effective technique for enhancing the reasoning capabilities of Large Language Models (LLMs) [33], with its underlying inference scaling laws having been empirically validated [34]. Reinforcement Learning (RL) represents a prominent methodology for implementing these scaling techniques [2]. The application of RL to LLMs has spurred two parallel lines of inquiry. The first focuses on the development of novel frameworks for post-training, such as VC-PPO [14], DAPO [15], and VAPO [16]. The second line of inquiry seeks to elucidate the fundamental impact of RL on LLM behavior, where researchers have explored diverse questions. For instance, Setlur et al. [35] demonstrated the conditional advantages of RL over Supervised Fine-Tuning (SFT). Yue et al. [36] empirically investigated whether RL can expand the knowledge boundaries of LLMs, while Shao et al. [37] examined the influence of reward modeling. Concurrently, Zhao et al. [38] posited that RL primarily serves to amplify behaviors acquired during pre-training. Although these studies provide valuable insights from different perspectives, the absence of a common theoretical formulation can lead to divergent conclusions.
5.3 Optimization and Generalization
Optimization in machine learning refers to the process of finding the minima of a loss function during model training. A prominent framework in this domain is Sharpness-Aware Minimization (SAM), which operates on the principle that flatter minima correlate with better generalization [39]. This phenomenon was later attributed to the inherent noise within the optimization process [40]. Building on this insight, subsequent works have proposed various techniques to enhance optimization, such as batch augmentation [41], extrapolation methods [42], and the introduction of a surrogate gap [43]. Generalization analysis evaluates a model’s ability to perform effectively on unseen data. Information theory provides a powerful theoretical foundation for this analysis. Key concepts include entropy, $H(X)$, which measures the information content of a random variable $X$, and mutual information, I(X;Y)=H(X)H(X|Y)I(X;Y)=H(X)-H(X|Y), which quantifies the information shared between two variables. Information-theoretic approaches have found wide-ranging applications in machine learning, from establishing bounds on the generalization capacity of deep learning models [44, 45] to enhancing model interpretability [46, 47, 48]. More recently, these tools have been employed to analyze synthetic data generation [49] and to quantify reasoning errors in LLMs [29].
6 Conclusion
In this paper, we addressed the critical misalignment between the token-level formulation of traditional reinforcement learning theories and the reasoning-level execution of Chain-of-Thought in Large Language Models. We introduced CoT-Space, a new theoretical framework that shifts the perspective from a discrete Markov Decision Process to a continuous optimization problem in a semantic reasoning space. This novel viewpoint models the generation of a reasoning chain as a trajectory towards a solution minimum in a reasoning loss landscape, providing a more intuitive and powerful lens for theoretical analysis.
Our analysis, conducted within the CoT-Space, validates the effectiveness of this framework. From an optimization perspective, we established an analogy between CoT length and the learning rate in classical machine learning, demonstrating that an optimal reasoning length corresponds to an optimal noise scale that promotes convergence to flatter, more generalizable solutions. Furthermore, from a risk analysis perspective, we decomposed the total reasoning error into empirical loss and generalization error. We proved that there exists a fundamental trade-off: shorter CoTs risk underfitting due to insufficient reasoning depth, while longer CoTs risk overfitting by increasing model complexity. This trade-off theoretically grounding the existence of an optimal CoT length $L_{opt}$, that minimizes total error.
Looking forward, the CoT-Space framework opens up several promising avenues for future research. First, it can be extended to analyze the impact of different RL algorithms and reward structures on the reasoning process, potentially leading to more principled algorithm design. Second, the insights from our risk analysis could inspire new regularization techniques aimed at explicitly controlling CoT complexity to prevent overfitting and improve generalization. Finally, this framework could be applied to other forms of complex reasoning beyond CoT, such as tree-based or graph-based thought processes, thereby advancing our theoretical understanding and practical ability to build more capable and reliable AI reasoning systems.
Appendix A Proofs
A.1 Proof of Theorem 2.6
Proof.
First, consider a single, fixed reasoning-level path $(\xi_{1},...,\xi_{L})$, where each step $\xi_{l}$ is realized by an average of $k$ tokens. By Assumption 2.5, the number of token-level realizations for this single path, $N_{\text{realize}}$, is bounded below:
$N_{\text{realize}}\geq\prod_{l=1}^{L}c^{k}=(c^{k})^{L}=c^{kL}.$ (4)
This shows that the number of token-level trajectories corresponding to even a single reasoning path grows exponentially with the total number of tokens, denoted as $K=kL$.
Next, we consider the total number of distinct, valid reasoning states, $|\mathcal{S}_{\text{reasoning}}^{(K)}|$. For complex problems, the number of distinct valid reasoning paths itself grows with $K$. As each additional reasoning step can potentially branch from multiple previous states, the number of reachable states grows super-linearly, often exponentially, with $K$. This implies $|\mathcal{S}_{\text{reasoning}}^{(K)}|\propto\alpha^{K}$ for some base $\alpha>1$.
From Definition 2.4, the reasoning state density $\rho(K)$ is thus:
$\rho(L)=\frac{|\mathcal{S}_{\text{reasoning}}^{(K)}|}{|\mathcal{V}_{\text{semantic}}|}\propto\frac{\alpha^{K}}{|\mathcal{V}_{\text{semantic}}|}.$ (5)
Assuming the semantic volume $\mathcal{V}_{\text{semantic}}$ for a given problem domain is fixed, the state density $\rho(K)$ grows exponentially with $K$.
Finally, we connect high density to the continuum property. In a $D$-dimensional space, the expected distance $\mathbb{E}[d(s_{i},s_{j})]$ between adjacent points $s_{i},s_{j}$ is inversely related to the density, approximately scaling as $\mathbb{E}[d(s_{i},s_{j})]\propto\rho(K)^{-1/D}$. Since $\rho(K)$ grows exponentially, it follows that:
$\lim_{K\to\infty}\mathbb{E}[d(s_{i},s_{j})]=0.$ (6)
When the expected distance between states approaches zero, it implies that for any state and any direction, another valid state can be found within an infinitesimally small neighborhood. This is the defining characteristic of a continuous space or manifold. Therefore, as the reasoning token amount $K$ increases, the discrete CoT-Space increasingly resembles a continuum, which validates the application of continuous optimization tools for its analysis. ∎
A.2 Proof of Theorem 3.1
Proof.
To bridge the discrete reasoning update with the continuous SDE model, we equate the variance induced by the noise term in both frameworks over a single reasoning step.
First, we analyze the noise in the discrete one-step reasoning update from Equation (1):
$\Delta s=-\frac{1}{L}\left(\frac{dC}{ds}+\left(\frac{d\hat{C}}{ds}-\frac{dC}{ds}\right)\right).$ (7)
The noise originates from the estimation error, which we define as $\alpha=\left(\frac{d\hat{C}}{ds}-\frac{dC}{ds}\right)$. We assume this error has a variance $\langle\alpha^{2}\rangle=F(s)$, where $F(s)$ is a function characterizing the magnitude of the reasoning loss estimation error at state $s$. The variance of the full noise term in the discrete update is therefore:
$\left\langle\left(-\frac{1}{L}\alpha\right)^{2}\right\rangle=\frac{1}{L^{2}}\langle\alpha^{2}\rangle=\frac{F(s)}{L^{2}}.$ (8)
Next, we analyze the noise from the SDE model proposed in Equation (2).
We assume the noise term $\eta(t)$ is white noise with the following statistical properties: $\langle\eta(t)\rangle=0$ and autocorrelation $\langle\eta(t)\eta(t^{\prime})\rangle=gF(s)\delta(t-t^{\prime})$, where $g$ is the noise scale. To find the variance over a discrete step, we integrate the SDE over a time interval $\Delta t=1/L$:
$\Delta s=\int_{0}^{1/L}\left(-\frac{dC}{ds}+\eta(t)\right)dt=-\frac{1}{L}\frac{dC}{ds}+\int_{0}^{1/L}\eta(t)dt.$ (9)
The variance of the noise component of this update is:
$\left\langle\left(\int_{0}^{1/L}\eta(t)dt\right)^{2}\right\rangle=\int_{0}^{1/L}\int_{0}^{1/L}\langle\eta(t)\eta(t^{\prime})\rangle dt^{\prime}dt=\int_{0}^{1/L}\int_{0}^{1/L}gF(s)\delta(t-t^{\prime})dt^{\prime}dt=\frac{gF(s)}{L}.$ (10)
By equating the variance from the discrete update (Eq. 8) and the SDE model (Eq. 10), we can solve for the noise scale $g$:
$\frac{F(s)}{L^{2}}=\frac{gF(s)}{L}.$ (11)
Solving for $g$, we find:
$g=\frac{1}{L}.$ (12)
This result shows that the noise scale $g$ in our CoT-Space SDE model is inversely proportional to the CoT length $L$. ∎
A.3 Proof and Analysis of Theorem 3.3
Proof.
Let’s denote the true reasoning loss of a policy $\pi$ as $R(\pi)=\mathbb{E}_{q\sim\mathcal{D}}[C(s_{q}^{\pi})]$ and the empirical reasoning loss as $\hat{R}_{n}(\pi)=\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi})$. The reasoning generalization error for a specific policy $\pi$ is $R(\pi)-\hat{R}_{n}(\pi)$. We aim to bound this gap.
The PAC-Bayes theorem provides a bound on the expected true loss for policies drawn from the posterior $Q$. A standard version of the theorem is often presented as follows.
Lemma A.1.
(PAC-Bayes generalization upper bound [18].) For any loss function bounded in $[0,1]$, for any prior $P$, with probability at least $1-\delta$ over the draw of the training set $\{q_{i}\}_{i=1}^{n}$, the following holds for all posterior distributions $Q$:
$\mathbb{E}_{\pi\sim Q}[R(\pi)]\leq\mathbb{E}_{\pi\sim Q}[\hat{R}_{n}(\pi)]+\sqrt{\frac{KL(Q||P)+\ln\frac{2n}{\delta}}{2n}},$
where $\operatorname{KL}(Q||P)=\mathbb{E}_{\pi\sim Q}[\ln\frac{Q(\pi)}{P(\pi)}]$ is the Kullback-Leibler (KL) divergence between the posterior and prior distributions.
Let’s adapt Lemma A.1 to our framework. We first need to normalize our reasoning loss $C(\cdot)$ to be in $[0,1]$. Assuming there is a maximum possible loss $C_{max}$, we can define a normalized loss $C^{\prime}(\cdot)=C(\cdot)/C_{max}$. The bound then applies to the normalized losses $R^{\prime}(\pi)$ and $\hat{R}^{\prime}_{n}(\pi)$. For simplicity of notation, let’s assume $C(\cdot)$ is already normalized in $[0,1]$. ∎
Information-Theoretic Interpretation
The bound has a clear information-theoretic meaning, which aligns with the goal of understanding reasoning phenomena:
  1. 1.
    Complexity Term ($\operatorname{KL}(Q||P)$): The KL divergence term measures the information gain in moving from our prior belief $P$ to our posterior belief $Q$ after seeing the data. It quantifies the complexity of the learned policy distribution.
    • If $Q$ is very different from $P$ (i.e., $\operatorname{KL}(Q||P)$ is large), it means we had to learn a great deal from the training data to find a good policy. This is analogous to overfitting or memorizing the training set. The bound becomes looser (larger), penalizing this complexity and suggesting a higher risk of poor generalization.
    • If $Q$ remains close to $P$ ($\operatorname{KL}(Q||P)$ is small), it implies the learned policy did not stray far from our initial bias towards simpler policies. The bound is tighter, suggesting better generalization.
  2. 2.
    Data Term ($\frac{1}{n}$): The bound depends on the number of training queries $n$. As $n$ increases, the term under the square root decreases, making the bound tighter. This confirms the intuition that having more data leads to better generalization and reduces the gap between training and test performance.
In conclusion, this PAC-Bayes bound formalizes the trade-off in learning a reasoning policy. To achieve good generalization, a policy $\pi$ must not only achieve a low empirical reasoning loss $\hat{R}_{n}(\pi)$ on the queries it has seen, but it must also remain "simple" in an information-theoretic sense, by not deviating too much from a reasonable prior. This framework provides a principled way to analyze and potentially control for phenomena like "overthinking" by viewing it as a form of overfitting, which would correspond to a large $\operatorname{KL}(Q||P)$ term.
A.4 Proof of Theorem 3.4
Proof.
The core idea is that a policy can only overfit to the training set $S=\{q_{1},\dots,q_{n}\}$ by encoding information about $S$ within its own parameters. The mutual information $I(\pi;S)$ precisely quantifies this amount of encoded information. A key result from learning theory provides the following bound on the expected generalization error:
Lemma A.2.
(Information-theoretical generalization upper bound [45].) With CoT-Space defined above, if the policy $\pi$ is trained on an i.i.d. drawn training set $S$ with size $n$, the following upper bound for the reasoning generalization error holds:
$|\mathbb{E}[R(\pi)-\hat{R}_{n}(\pi)]|\leq\sqrt{\frac{C_{max}^{2}\cdot I(\pi;S)}{2n}},$
where $C_{max}$ is the maximum possible value of the reasoning loss $C(\cdot)$, and the expectation is over the random draw of the training set $S$ and the (potentially stochastic) output $\pi$ of the learning algorithm.
Lemma A.2 transforms the problem of bounding generalization error into the problem of bounding the mutual information $I(\pi;S)$. To make this bound meaningful for our framework, we must connect $I(\pi;S)$ to the structural properties of the generated Chain-of-Thought.
We then turn to bound mutual information by CoT complexity. The policy $\pi$ expresses the information it has learned from $S$ by generating a CoT, which is a sequence of reasoning steps $\xi=(\xi_{1},\xi_{2},\dots,\xi_{L})$. Each step $\xi_{i}$ is, in turn, a sequence of tokens. The total information capacity of the policy’s output is therefore determined not just by the number of steps, but by the total number of tokens generated.
Let us define:
  • $L$: The number of reasoning steps in a CoT.
  • $|\xi_{i}|$: The number of tokens in the $i$-th reasoning step.
  • $\sum_{i=1}^{L}|\xi_{i}|$: The total number of tokens in the CoT.
The mutual information $I(\pi;S)$ is upper-bounded by the entropy of the policy’s output, which represents its description length or information capacity. The total number of token-level decisions the policy makes determines this capacity. Let $\mathbb{E}[L]$ be the expected number of reasoning steps (reasoning depth) and $\mathbb{E}[|\xi|]$ be the expected number of tokens per step (step verbosity). The expected total number of tokens is $\mathbb{E}[L]\cdot\mathbb{E}[|\xi|]$.
If we assume the policy generates tokens from a dictionary $\mathcal{A}$, the information required to specify one token is at most $\log|\mathcal{A}|$. Therefore, we can establish the following upper bound on the mutual information.
Lemma A.3.
In LLM reasoning scenarios, if the policy $\pi$ is trained on a training set $S$, the mutual information $I(\pi;S)$ is upper bounded as follows:
$I(\pi;S)\leq\mathbb{E}_{S,\pi}\left[\sum_{i=1}^{L}|\xi_{i}|\right]\cdot\log|\mathcal{A}|=\left(\mathbb{E}[L]\cdot\mathbb{E}[|\xi|]\right)\cdot\log|\mathcal{A}|.$
Lemma A.3 formalizes the intuition that the information a policy learns from data cannot exceed the information content it is capable of expressing in its output. Together with Lemma A.2, this finishes the proof. ∎
A.5 Proof of Theorem 3.7
Proof.
We begin with its definition:
$\hat{R}_{n}(\pi)=\frac{1}{n}\sum_{i=1}^{n}C(s_{q_{i}}^{\pi}).$ (13)
We can partition the sum over the training set into two disjoint subsets based on whether the generated CoT length was sufficient:
  • The set of failed queries: $S_{fail}=\{q_{i}\in S\mid L_{i}<L^{*}(q_{i})\}$.
  • The set of potentially successful queries: $S_{succ}=\{q_{i}\in S\mid L_{i}\geq L^{*}(q_{i})\}$.
Let $n_{fail}=|S_{fail}|$. We can now rewrite the empirical loss as:
$\hat{R}_{n}(\pi)=\frac{1}{n}\left(\sum_{q_{i}\in S_{fail}}C(s_{q_{i}}^{\pi})+\sum_{q_{i}\in S_{succ}}C(s_{q_{i}}^{\pi})\right).$ (14)
Applying our failure assumption, we know that for every $q_{i}\in S_{fail}$, $C(s_{q_{i}}^{\pi})\geq C_{fail}$. For queries in $S_{succ}$, the loss is at least 0. This allows us to establish a lower bound:
$\hat{R}_{n}(\pi)\geq\frac{1}{n}\left(\sum_{q_{i}\in S_{fail}}C_{fail}+\sum_{q_{i}\in S_{succ}}0\right)=\frac{n_{fail}\cdot C_{fail}}{n}.$ (15)
We can define the empirical failure rate of the policy $\pi$ on the training set as $P_{\pi}(L<L^{*})=n_{fail}/n$. This leads to our final lower bound for the empirical loss:
$\hat{R}_{n}(\pi)\geq P_{\pi}(L<L^{*})\cdot C_{fail}.$ (16)
This finishes the proof. ∎
Appendix B More Experiments
Figure 5: Supplementary validation of the relationship between task difficulty and optimal CoT length ($L_{\text{opt}}$) on additional models. The plots illustrate that the converged mean response length ($L_{\text{opt}}$) consistently increases with task difficulty across three different models: (a) Qwen2.5-3B-Base, (b) Qwen2.5-1.5B-Base, and (c) Qwen2.5-0.5B-Base. These results reinforce the findings presented in Figure 4(a) and demonstrate the robustness of Remark 1 across models of varying capacities.
To further strengthen the empirical validation of Remark 1, we conducted supplementary experiments to verify the relationship between task difficulty and the optimal CoT length ($L_{\text{opt}}$) across models of varying capacities. The experimental protocol is identical to that described in Subsection 4.2, but here we replicate the analysis on three additional models from the Qwen2.5 series: Qwen2.5-3B-Base, Qwen2.5-1.5B-Base, and Qwen2.5-0.5B-Base.
The results, presented in Figure 5, consistently corroborate the findings from our main experiment (Figure 4(a)). Across all three smaller models, we observe the same clear, monotonic trend: the converged optimal CoT length ($L_{\text{opt}}$) increases systematically as the difficulty of the task escalates from GSM8K to the higher levels of the MATH dataset. This demonstrates that the principle outlined in Remark 1, that more complex problems necessitate a longer reasoning process to avoid underfitting is not specific to a single high-capacity model but holds true as a general principle across models of different scales.